# Running in circles: is practical application feasible for data fission and data thinning in post-clustering differential analysis?

### v1

@@ 4 Discussion
In the Gaussian framework, we theoretically quantified the relationship between the relative bias in variance estimation and the associated Type I error in post-clustering $t$-tests. In practice, our simulation results suggest that a small relative bias can be acceptable while still achieving effective Type I error control. These first results pave the way for defining a principled approach to tuning the hyperparmater $tau$ in data fission and data thining for post-clustering inference to optimize statistical power.
@@
As a solution to avoid the need for prior knowledge of component-specific variance parameters in the Gaussian mixture setting, we propose a heteroscedastic model with individual variances, that can be estimated through a non-parametric weighted estimator. This approach aligns more closely with the distributional assumptions made by data fission and thinning. However, the performance of this non-parametric approach relies heavily on the choice of its bandwidth. The best bandwidth would be the one capturing only the observations originating from a same components, but it would again require knowledge of the true mixture components whereas their estimation is part of the clustering first step of the method.
@@
Finally, we demonstrated that the results derived the Gaussian distribution context are readily applicable to the negative binomial distribution commonly employed in modeling RNA-seq data. Specifically, we illustrate on real data that overdispersion is also component-specific. Therefore, without knowing the true cluster structure data, data thinning can not ensure the needed independence between clustering and differential testing.
@@
In practice, the application of data fission or data thinning for post-clustering inference appears to be akin to a self-referential loop generating circular reasoning. While introduced as a solution for addressing challenges in post-clustering inference, all strategies that could theoretically ensure independence between the two stages of the analysis ultimately rely on knowledge of the true, but unknown, clusters. Despite its conceptual appeal, the practical utility of these methods for post-clustering inference is remains limited to extreme cases with extremely high signal vs. noise ratios, emphasizing the need for alternative methodologies that can navigate the complexities of unknown class structures more effectively.
@@
All codes and data needed to reproduce the results presented here are openly accessible from Zenodo with DOI 10.5281/zenodo.11207777 (Hivert, Agniel, Thiébaut & Hejblum, 2024).
@@ 5 Acknowledgements
BH is supported partly by the Digital Public Health Graduate’s school, funded by the PIA $3$ (Investments for the Future – Project reference: 17-EURE-0019). The work was supported through the DESTRIER Inria Associate-Team from the Inria@SiliconValley program (analytical code: DRI-012215), and by the CARE project funded from the Innovative Medicines Initiative 2 Joint Undertaking (JU) under grant agreement no IMI2-101005077. The JU receives support from the European Union’s Horizon $2020$ research and innovation programme and EFPIA and Bill &\&& Melinda Gates Foundation, Global Health Drug Discovery Institute, University of Dundee. This work has received funding from the European Union’s Horizon $2020$ research and innovation programme under EHVA grant agreement no H2020-681032. This study was carried out in the framework of the University of Bordeaux’s France 2030 program / RRI PHDS. Computer time for this study was provided by the computing facilities MCIA (Mésocentre de Calcul Intensif Aquitain) of the Université de Bordeaux and of the Université de Pau et des Pays de l’Adour.
Appendix A Independence proof of the Gaussian process
Let $\boldsymbol{X}\sim\mathcal{N}\left(\boldsymbol{\mu},\boldsymbol{\Sigma}_{p\times p}\right)$. Considering the fission process for Gaussian data described in Table 1, we can decompose $\boldsymbol{X}$ into two new random variables $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$ using a new random variable $\boldsymbol{Z}\sim\mathcal{N}\left(\boldsymbol{0},\boldsymbol{\Sigma}\right)$. It follows form this decomposition that:
@@ Appendix A Independence proof of the Gaussian process
Appendix B Impact of covariance estimation under the independence between $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$
@@
Appendix C Derivation of $t$-test statistic under the null of no-cluster in the univariate data fission post-clustering setting
The fission process, as outlined in Table 1, gives:
$X^{(1)}\sim\mathcal{N}(\mu,\sigma^{2}+\tau^{2}\widehat{\sigma^{2}})\quad\text{and}\quad X^{(2)}\sim\mathcal{N}\left(\mu,\sigma^{2}+\frac{1}{\tau^{2}}\widehat{\sigma^{2}}\right)$
After applying k-means clustering to the $n$ realizations of $X^{(1)}$, two clusters $\widehat{C}_{1}$ and $\widehat{C}_{2}$ are obtained. As $X^{(1)}$ follows a Gaussian distribution, the k-means clustering splits it into two parts $X^{(1)}|C_{k}$, $k=1,2$, with the same size $n/2$ and the same variance, forming two half-Gaussian distributions. It follows from the half-Gaussian distribution that:
$\mathbb{E}[X^{(1)}|C_{k}]=\pm\sqrt{\frac{2\left(\sigma^{2}+\tau^{2}\widehat{\sigma^{2}}\right)}{\pi}}$ (5)
where the sign depends on whether $\hat{C}_{k}$ represents the truncation on the right or left side of $\mu$. Using (5), the mean differences between $\widehat{C}_{1}$ and $\widehat{C}_{2}$ is then given by $\hat{\delta}_{X^{(1)}}=\left|\mathbb{E}\left[X^{(1)}|\widehat{C}_{2}\right]-\mathbb{E}\left[X^{(1)}|\widehat{C}_{1}\right]\right|=2\sqrt{\frac{2}{\pi}\left(\sigma^{2}+\tau^{2}\widehat{\sigma^{2}}\right)}$. Due to the k-means clustering, the two estimated clusters share the same intra-cluster variance. Following the same half-Gaussian arguments, this shared intra-cluster variance in $X^{(1)}$ is given by:
$s_{X^{(1)}}^{2}=(\sigma^{2}+\tau^{2}\widehat{\sigma^{2}})\left(1-\frac{2}{\pi}\right)$
In the post-clustering data fission setting, hypothesis testing is applied on the realizations of $X^{(2)}$ using $\widehat{C}_{1}$ and $\widehat{C}_{2}$. Our focus is on the mean differences $\hat{\delta}_{X^{(2)}}$ between these two clusters and their shared intra-cluster variances $s^{2}(X^{(2)})$ on $X^{(2)}$. These values depend on the ability to reproduce the clustering results on $X^{(2)}$ which is quantified by $\operatorname{Cor}\left(X^{(1)},X^{(2)}\right)$. Assuming the same clustering for $X^{(2)}$ as for $X^{(1)}$ implies the same truncation on $X^{(2)}$. In this case, because $X^{(2)}\sim\mathcal{N}\left(\mu,\sigma^{2}+\frac{1}{\tau^{2}}\widehat{\sigma^{2}}\right)$, we have:
$\hat{\delta}_{X^{(2)}}=2\sqrt{\frac{2}{\pi}\left(\sigma^{2}+\frac{1}{\tau^{2}}\widehat{\sigma^{2}}\right)}$
However, considering that the cluster analysis is applied first to $X^{(1)}$ and then transferred to $X^{(2)}$, we need to account for the correlation between $X^{(1)}$ and $X^{(2)}$, leading to:
$\hat{\delta}_{X^{(2)}}$ $=2\sqrt{\frac{2}{\pi}\left(\sigma^{2}+\frac{1}{\tau^{2}}\widehat{\sigma^{2}}\right)}\times\left|\operatorname{Cor}\left(X^{(1)},X^{(2)}\right)\right|$
$=2\sqrt{\frac{2}{\pi}\left(\sigma^{2}+\frac{1}{\tau^{2}}\widehat{\sigma^{2}}\right)}\times\left|\frac{(\sigma^{2}-\widehat{\sigma^{2}})}{\sqrt{(\sigma^{2}+\tau^{2}\widehat{\sigma^{2}})\times(\sigma^{2}+\frac{1}{\tau^{2}}\widehat{\sigma^{2}})}}\right|$
This mean difference $\hat{\delta}_{X^{(2)}}$ can be interpreted as a trade-off between how easily $\widehat{C}_{1}$ and $\widehat{C}_{2}$ can be reproduced on $X^{(2)}$ and the mean differences that would be observed if they were perfectly reproduced on $X^{(2)}$. For the shared intra-cluster variance, $s^{2}(X^{(2)})$ should approximate $(\sigma^{2}+\frac{1}{\tau}^{2}\widehat{\sigma^{2}})(1-\frac{2}{\pi})$ when $\left|\operatorname{Cor}\left(X^{(1)},X^{(2)}\right)\right|$ is close to $1$ (the scenario where we accurately replicate $\widehat{C}_{1}$ and $\widehat{C}_{2}$ on $X^{(2)}$). In this situation, as the two estimated clusters are fully preserved, their variances precisely follow the half-Gaussian variance on $X^{(2)}$. Conversely, when $\left|\operatorname{Cor}\left(X^{(1)},X^{(2)}\right)\right|$ is close to 00 (indicating the inability to reproduce the clustering results on $X^{(2)}$ due to its independence with $X^{(1)}$), $s^{2}(X^{(2)})$ should approximate $\sigma^{2}+\frac{1}{\tau^{2}}\widehat{\sigma^{2}}$, representing the entire variance of $X^{(2)}$. This is a consequence of the independence between $X^{(1)}$ and $X^{(2)}$ that leads to random clusters on $X^{(2)}$. Thus, these two considerations give:
$s^{2}(X^{(2)})=\left(\sigma^{2}+\frac{1}{\tau^{2}}\widehat{\sigma^{2}}\right)\left(1-\frac{2}{\pi}\right)\operatorname{Cor}\left(X^{(1)},X^{(2)}\right)^{2}+\left(\sigma^{2}+\frac{1}{\tau^{2}}\widehat{\sigma^{2}}\right)\left(1-\left(\operatorname{Cor}\left(X^{(1)},X^{(2)}\right)\right)^{2}\right)$ (6)
In the post-clustering inference setting, we are interested in testing a mean difference between the two estimated clusters $\widehat{C}_{1}$ and $\widehat{C}_{2}$. So, the null hypothesis we are interested in is:
$\mathcal{H}_{0}:\mu_{C_{1}}=\mu_{C_{2}}\qquad\mbox{vs}\qquad\mathcal{H}_{1}:\mu_{C_{1}}\neq\mu_{C_{2}}$
Under the null, $T_{\widehat{C}_{1},\widehat{C}_{2}}=\frac{\overline{X^{(2)}_{\widehat{C}_{1}}}-\overline{X^{(2)}_{\widehat{C}_{2}}}}{\sqrt{\frac{\hat{s}^{2}_{\widehat{C}_{1}}}{n_{1}}+\frac{\hat{s}^{2}_{\widehat{C}_{2}}}{n_{2}}}}$ where $\overline{X^{(2)}_{C_{k}}}=\frac{1}{n_{k}}\sum\limits_{i\in C_{k}}x_{i}^{(2)}$.
We assumed that the $2-$means clustering results in clusters such that $n_{1}=n_{2}=\frac{n}{2}$. Moreover, using the half Gaussian argument, we supposed that $\widehat{C}_{1}$ and $\widehat{C}_{2}$ shares the same intra-cluster variances given by $s^{2}(X^{(2)})$. So,
$T_{\widehat{C}_{1},\widehat{C}_{2}}$ $=\frac{\overline{X^{(2)}_{\widehat{C}_{1}}}-\overline{X^{(2)}_{\widehat{C}_{2}}}}{\sqrt{\frac{s^{2}(X^{(2)})}{\frac{n}{2}}+\frac{s^{2}(X^{(2)})}{\frac{n}{2}}}}$
$=\frac{\overline{X^{(2)}_{\widehat{C}_{1}}}-\overline{X^{(2)}_{\widehat{C}_{2}}}}{\sqrt{\frac{4s^{2}(X^{(2)})}{n}}}$
We want to compute the null distribution of $T_{\widehat{C}_{1},\widehat{C}_{2}}$. Using the Central Limit theorem, we know that:
$\overline{X^{(2)}}_{\widehat{C}_{1}}-\overline{X^{(2)}}_{\widehat{C}_{2}}\sim\mathcal{N}\left(\underbrace{\mathbb{E}[X^{(2)}|\widehat{C}_{1}]-\mathbb{E}[X^{(2)}|\widehat{C}_{2}]}_{\hat{\delta}_{X^{(2)}}},\operatorname{\mathbb{V}ar}\left(\overline{X^{(2)}_{\widehat{C}_{1}}}-\overline{X^{(2)}_{\widehat{C}_{2}}}\right)\right)$
But,
$\operatorname{\mathbb{V}ar}\left(\overline{X^{(2)}_{\widehat{C}_{1}}}-\overline{X^{(2)}_{\widehat{C}_{2}}}\right)$ $=\operatorname{\mathbb{V}ar}\left(\overline{X^{(2)}_{\widehat{C}_{1}}}\right)+\operatorname{\mathbb{V}ar}\left(\overline{X^{(2)}_{\widehat{C}_{2}}}\right)\qquad\mbox{by independence}$
@@
$=\frac{2}{\left(\frac{n}{2}\right)^{2}}\sum\limits_{i\in\widehat{C}_{1}}\operatorname{\mathbb{V}ar}\left(X^{(2)}_{i}|C_{1}\right)$
$=\frac{8}{n^{2}}\frac{n}{2}\operatorname{\mathbb{V}ar}\left(X^{(2)}_{1}|C_{1}\right)$
$=\frac{4}{n}\operatorname{\mathbb{V}ar}\left(X^{(2)}_{1}|C_{1}\right)$
$=\frac{4}{n}s^{2}(X^{(2)})$
So,
$\overline{X^{(2)}}_{\widehat{C}_{1}}-\overline{X^{(2)}}_{\widehat{C}_{2}}\sim\mathcal{N}\left(\widehat{\delta}_{X^{(2)}},\frac{4}{n}s^{2}(X^{(2)})\right)$
Finally, we have:
$\mathbb{E}\left[T_{\widehat{C}_{1},\widehat{C}_{2}}\right]$ $=\mathbb{E}\left[\frac{\overline{X^{(2)}_{\widehat{C}_{1}}}-\overline{X^{(2)}_{\widehat{C}_{2}}}}{\sqrt{\frac{4s^{2}(X^{(2)})}{n}}}\right]$
$=\frac{1}{\sqrt{\frac{4s^{2}(X^{(2)})}{n}}}\mathbb{E}\left[\overline{X^{(2)}_{\widehat{C}_{1}}}-\overline{X^{(2)}_{\widehat{C}_{2}}}\right]$
$=\frac{\widehat{\delta}_{X^{(2)}}}{\sqrt{\frac{4s^{2}(X^{(2)})}{n}}}$
and:
$\operatorname{\mathbb{V}ar}\left(T_{\widehat{C}_{1},\widehat{C}_{2}}\right)$ $=\operatorname{\mathbb{V}ar}\left(\frac{\overline{X^{(2)}_{\widehat{C}_{1}}}-\overline{X^{(2)}_{\widehat{C}_{2}}}}{\sqrt{\frac{4s^{2}(X^{(2)})}{n}}}\right)$
$=\frac{1}{\frac{4s^{2}(X^{(2)})}{n}}\operatorname{\mathbb{V}ar}\left(\overline{X^{(2)}_{\widehat{C}_{1}}}-\overline{X^{(2)}_{\widehat{C}_{2}}}\right)$
$=\frac{1}{\frac{4s^{2}(X^{(2)})}{n}}\frac{4s^{2}(X^{(2)})}{n}$
$=1$
So,
$T_{\widehat{C}_{1},\widehat{C}_{2}}\mathop{\sim}\limits^{\mathcal{H}_{0}}\mathcal{N}\left(\frac{\widehat{\delta}_{X^{(2)}}}{\sqrt{\frac{4s^{2}(X^{(2)})}{n}}},1\right)$
Since we have, $\widehat{\delta}_{X^{(2)}}=2\sqrt{\frac{2}{\pi}\left(\sigma^{2}+\frac{1}{\tau^{2}}\widehat{\sigma^{2}}\right)}\times\left|\operatorname{Cor}\left(X^{(1)},X^{(2)}\right)\right|$ and $s^{2}(X^{(2)})=(\sigma^{2}+\frac{1}{\tau^{2}}\widehat{\sigma^{2}})(1-\frac{2}{\pi})\left(\operatorname{Cor}\left(X^{(1)},X^{(2)}\right)\right)^{2}+(\sigma^{2}+\frac{1}{\tau^{2}}\widehat{\sigma^{2}})\left(1-\left(\operatorname{Cor}\left(X^{(1)},X^{(2)}\right)\right)^{2}\right)$, which can be simplified as $\left(\sigma^{2}+\frac{1}{\tau^{2}}\widehat{\sigma^{2}}\right)\left(1-\frac{2}{\pi}\operatorname{Cor}\left(X^{(1)},X^{(2)}\right)^{2}\right)$, then:
$\frac{\widehat{\delta}_{X^{(2)}}}{\sqrt{\frac{4s^{2}(X^{(2)})}{n}}}$ $=\frac{2\sqrt{\frac{2}{\pi}\left(\sigma^{2}+\frac{1}{\tau^{2}}\widehat{\sigma^{2}}\right)}\times\left|\operatorname{Cor}\left(X^{(1)},X^{(2)}\right)\right|}{\sqrt{\frac{4\left(\sigma^{2}+\frac{1}{\tau^{2}}\widehat{\sigma^{2}}\right)\left(1-\frac{2}{\pi}\operatorname{Cor}\left(X^{(1)},X^{(2)}\right)^{2}\right)}{n}}}$
@@
$=\frac{2\sqrt{\frac{2}{\pi}}\sqrt{\left(\sigma^{2}+\frac{1}{\tau^{2}}\widehat{\sigma^{2}}\right)}\sqrt{\operatorname{Cor}\left(X^{(1)},X^{(2)}\right)^{2}}}{\frac{2\sqrt{\left(\sigma^{2}+\frac{1}{\tau^{2}}\widehat{\sigma^{2}}\right)}\sqrt{1-\frac{2}{\pi}\operatorname{Cor}\left(X^{(1)},X^{(2)}\right)^{2}}}{\sqrt{n}}}$

### v2

@@ 4 Discussion
In the Gaussian framework, we theoretically quantify the relationship between the relative bias in variance estimation and the associated Type I error in post-clustering $t$-tests. In practice, our simulation results suggest that a small relative bias can be acceptable while still achieving effective Type I error control. These first results pave the way for defining a principled approach to tuning the hyperparmater $\tau$ in data fission and data thining for post-clustering inference to optimize statistical power.
@@
As a solution to avoid the need for prior knowledge of component-specific variance parameters in the Gaussian mixture setting, we propose a heteroscedastic model with individual variances, that can be replaced by a plug-in estimate (such as a non-parametric estimator of the local variance). This approach aligns more closely with the distributional assumptions made by data fission and thinning. However, the performance of our non-parametric approach relies heavily on the choice of its bandwidth. The best bandwidth would be the one capturing only the observations originating from the same component, but it would again require knowledge of the true mixture components in spite of their estimation being part of the first clustering step of the method. Consequently, while this heteroscedastic model better fits the parametric assumptions of data fission and thinning, leveraging an accurate enough plug-in estimator remains challenging. This underscores the difficulty of adapting these methods to mixture distributions. We show that when the signal to noise ratio is extremely favorable for the clustering, with very well-separated components, this approach unlocks the use of data-fission for post-clustering differential testing of clustered data. In practice though, one can question the need for data fission in such cases as the uncertainty regarding the clustering is actually very small when $\delta/\sigma\geq 10$. We have also investigated iteratively fissioning the data and updating the variance plug-in estimate, but this heuristic strategy requires even larger separation of the components to effectively work.
@@
Finally, we demonstrate that the results derived from the Gaussian distribution context are readily applicable to the negative binomial distribution commonly employed in modeling RNA-seq data. Specifically, we illustrate on real data that overdispersion is also component-specific. Therefore, without knowing the true cluster structure in the data, data thinning cannot ensure the needed independence between clustering and differential testing.
@@
In practice, the application of data fission or data thinning for post-clustering inference appears to be akin to a self-referential loop generating circular reasoning. While introduced as a solution for addressing challenges in post-clustering inference, all strategies that could theoretically ensure independence between the two stages of the analysis ultimately rely on knowledge of the true, but unknown, clusters. Despite its conceptual appeal, the practical utility of these methods for post-clustering inference remains limited to extreme cases with extremely high signal vs. noise ratios, emphasizing the need for alternative methodologies that can navigate the complexities of unknown class structures more effectively.
@@
All codes and data needed to reproduce the results presented here are openly accessible from Zenodo with DOI 10.5281/zenodo.11207777 (Hivert et al., 2024b, ).
@@ 5 Acknowledgements
BH is supported partly by the Digital Public Health Graduate’s school, funded by the PIA $3$ (Investments for the Future – Project reference: 17-EURE-0019). The work was supported through the DESTRIER Inria Associate-Team from the Inria@SiliconValley program (analytical code: DRI-012215), and by the CARE project funded from the Innovative Medicines Initiative 2 Joint Undertaking (JU) under grant agreement no IMI2-101005077. The JU receives support from the European Union’s Horizon $2020$ research and innovation programme and EFPIA and Bill &\&& Melinda Gates Foundation, Global Health Drug Discovery Institute, University of Dundee. This work has received funding from the European Union’s Horizon $2020$ research and innovation programme under EHVA grant agreement no H2020-681032. This study was carried out in the framework of the University of Bordeaux’s France 2030 program / RRI PHDS. This work benefited from State aid managed by the Agence Nationale de la Recherche under the France 2030 program, reference AI4scMED ANR-22-PESN-0002. Computer time for this study was provided by the computing facilities MCIA (Mésocentre de Calcul Intensif Aquitain) of the Université de Bordeaux and of the Université de Pau et des Pays de l’Adour.
Appendix S1 Independence proof of the Gaussian process
Let $\boldsymbol{X}\sim\mathcal{N}\left(\boldsymbol{\mu},\boldsymbol{\Sigma}_{p\times p}\right)$. Considering the fission process for Gaussian data described in Table $1$, we can decompose $\boldsymbol{X}$ into two new random variables $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$ using a new random variable $\boldsymbol{Z}\sim\mathcal{N}\left(\boldsymbol{0},\boldsymbol{\Sigma}\right)$. It follows form this decomposition that:
@@
Appendix S2 Impact of covariance estimation under the independence between $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$
@@
Appendix S3 Derivation of $t$-test statistic under the null of no-cluster in the univariate data fission post-clustering setting
@@
$X^{(1)}_{i}=X_{i}+\tau Z_{i}\qquad\text{and}\qquad X^{(2)}_{i}=X_{i}-\frac{1}{\tau}Z_{i}$
We can immediately deduce the marginal distributions of $X^{(1)}_{i}$ and $X^{(2)}_{i}$ for all $i=1,\dots,n$, thanks to the independence between $X_{i}$ and $Z_{i}$:
$X^{(1)}_{i}\sim\mathcal{N}\left(\mu,\sigma^{2}+\tau^{2}b^{2}\right)\qquad\text{and}\qquad X^{(2)}_{i}\sim\mathcal{N}\left(\mu,\sigma^{2}+\frac{1}{\tau^{2}}b^{2}\right)$ (5)
We denote $\sigma^{2}_{X^{(1)}}=\sigma^{2}+\tau^{2}b^{2}$ and $\sigma^{2}_{X^{(2)}}=\sigma^{2}+\frac{1}{\tau^{2}}b^{2}$ as the respective variances of $X^{(1)}_{i}$ and $X^{(2)}_{i}$ above.
In the context of data fission to address the challenges of post-clustering inference, a clustering algorithm is applied to the observations of $\boldsymbol{X}^{(1)}$. Without loss of generality, we assume that the clustering algorithm applied to the realizations separates $X^{(1)}_{1},\dots,X^{(1)}_{n}$ into two clusters $C_{1}$ and $C_{2}$ around $\mu$ (which is typically the case with the $k$-means algorithm or with a two-component Gaussian mixture model with homogeneous variance when $n$ is sufficiently large). We also assume that these two clusters have the same size and the same variance.
Thus, the clusters $C_{1}$ and $C_{2}$ can be expressed as:
$C_{1}=\left\{i=1,\dots,n:X^{(1)}_{i}>\mu\right\}\qquad\text{and}\qquad C_{2}=\left\{i=1,\dots,n:X^{(1)}_{i}\leq\mu\right\}$
We can then derive the conditional distributions of $X^{(1)}_{i}|C_{1}$ and $X^{(1)}_{i}|C_{2}$. Indeed, $\mathbb{P}(X^{(1)}_{i}=x|C_{1})=\mathbb{P}(X^{(1)}_{i}=x|X^{(1)}_{i}>\mu)$ with $X^{(1)}_{i}\sim\mathcal{N}\left(\mu,\sigma^{2}_{X^{(1)}}\right)$. Therefore, $X^{(1)}_{i}|X^{(1)}_{i}>\mu$ follows a half-normal distribution, and for all $i=1,\dots,n$:
@@
$\mathbb{E}\left[X^{(1)}_{i}|X^{(1)}_{i}>\mu\right]=\mu+\sqrt{\frac{2\sigma^{2}_{X^{(1)}}}{\pi}}\quad\text{and}\quad\operatorname{Var}\left(X^{(1)}_{i}|X^{(1)}_{i}>\mu\right)=\left(1-\frac{2}{\pi}\right)\sigma^{2}_{X^{(1)}}$
Since the cluster $C_{2}=\left\{i=1,\dots,n:X^{(1)}_{i}\leq\mu\right\}$ simply represents the cluster on the other side of the mean $\mu$, we similarly have: